How long does it take for the current in an RL circuit with R = 130 W and L = 63 mH to reach half its final value?
Answer #1

R = 130 O

L = 63*10^-3 H

i = i0/2

t =?

we know that

i = i0 [1 - e-Rt/L ]

i0/2 = i0 [1 - e-Rt/L ]

e-Rt/L= 1/2

eRt/L = 2

Taking ln on both sides,we get

Rt/L = ln 2

(130*t)/(63*10^-3) = 0.693

t = 3.358*10-4 s = 0.3358 ms